# Swarming in the Wild: A Distributed Communication-less Lloyd-based Algorithm dealing with Uncertainties

### v1

@@ Abstract
In this work, we present a distributed algorithm for swarming in complex environments that operates with no communication, no a priori information about the environment, and using only onboard sensing and computation capabilities. We provide sufficient conditions to guarantee that each robot reaches its goal region in a finite time, avoiding collisions with obstacles and other robots without exceeding a desired maximum distance from a predefined set of neighbors (flocking constraint). In addition, we show how the proposed algorithm can deal with tracking errors and onboard sensing errors without violating safety and proximity constraints, still providing the conditions for having convergence towards the goal region. To validate the approach, we provide experiments in the field. We tested our algorithm in GNSS-denied environments i.e., a dense forest, where fully autonomous aerial robots swarmed safely to the desired destinations, by relying only on onboard sensors, i.e., without a communication network. This work marks the initial deployment of a fully distributed system where there is no communication between the robots, nor reliance on any global localization system, which at the same time it ensures safety and convergence towards the goal within such complex environments.
@@ I Introduction
Over the past few decades, numerous applications have emerged for multi-robot systems, exhibiting their undisputed benefits in diverse fields. These applications span domains such as logistics, surveillance, environment monitoring, disaster response, search and rescue, and agriculture among others, demonstrating their versatility. For many of these applications, having a swarm of robots able to flock together towards a desired location can be highly beneficial, due to redundancies and parallelization. The definition of flocking is provided by the seminal work of Reynolds [1], which defines it as a non-colliding aggregate motion. Depending on the environments where robots operate, specific factors like the reliability or availability of direct communication between robots may be compromised.
In this work, we propose a Lloyd-based algorithm that relies only on on-board sensors. We aim to synthesize a distributed control algorithm, where each robot safely reaches the desired position, maintaining proximity between robots in the team. In particular, we provide sufficient conditions to guarantee safety, proximity maintenance and convergence towards the goal in cluttered environments by accounting for bounded tracking errors and uncertainties in the estimation of neighbors and obstacles positions.
Our algorithm takes inspiration from nature; in fact, just like flock of birds or a school of fish, the robots do not need a leader and each robot makes its own decisions based on relative information, without relying on explicit communication. We validate our approach through an extensive simulation and experimental campaign in the field. In Figure 1, we illustrate a snapshot during a mission in the forest. This work represents the first deployment of a fully distributed system with absence of any communication between the robots and any global localization system, which also provides conditions for safety and convergence towards the goal in such a complex environments.
Figure 1: Deployment of the proposed algorithm in a real-world scenario with $3$ UAVs navigating in a dense forest.
@@ I-A Related works
Prior literature provides a considerable amount of algorithms for distributed motion coordination of multiple robots, but only a few of them can manage to work in a cluttered environment by also imposing the flocking constraint. Among them, a smaller subset is devoted to applications in the field without relying on direct communication between the robots.
To position our algorithm within the state of the art, we considered the following features:
  • 1.
    Distributed control: each robot makes decisions on the basis of local information, which is obtained only by relying on onboard sensors;
  • 2.
    Asynchronous control: each robot makes decisions independently, in asynchronous fashion;
  • 3.
    Communication independence: the robots do not need to communicate, since the algorithm does not rely on unobservable information, e.g., future trajectories, velocities, and accelerations;
  • 4.
    Safety: the system is provably safe, i.e., given the upper bounds on uncertainties (see item 7), collision and obstacle avoidance is guaranteed at any time;
  • 5.
    Convergence: the sufficient conditions for the robots to converge to their goal locations are provided (i.e., absence of deadlock and live-lock);
  • 6.
    Proximity maintenance: the algorithm allows to set hard constraints on the maximum distance between robots;
  • 7.
    Uncertainties: the system is provably safe and respects proximity constraints even in the presence of bounded tracking errors and/or bounded measurement uncertainties;
  • 8.
    Cluttered and highly dense environments: the algorithm can safely manage the presence of static obstacles in the environment. It can also handle highly dense cluttered environments, such as a dense forest.
  • 9.
    Onboard localization: the algorithm is tested by relying on onboard localization in unknown environments.
  • 10.
    Experiments: the proposed algorithm is tested on real robotic platforms in unstructured environments;
  • 11.
    Open source code: the algorithm is open source on Github.
@@
Table I: List of state of the art algorithms. We indicate with “$\times$” if the algorithm does not implement the feature, “✓” if it does, and “o” if it does only partially.
In Table I, we report an overview of the most closely related algorithms in the literature. We mark with “$\times$” if the algorithm does not implement the feature, “✓” if it does, and finally, “o” if it does but only partially. As shown in Table I, our algorithm is unmatched in fulfilling all the desired features of interest. In fact, despite [5, 4] proposing valid solutions to the problem, their main limitations are the requirement of a trajectory broadcasting network, the absence of both convergence analysis and a rigorous account of tracking errors and uncertainties. Similarly, in [6] the robots need to broadcast information such as poses, velocities, poses covariance, and global extrinsic transformations. On the other hand, [7, 8], do not rely on a communication network, but they exhibit degraded performance in crowded scenarios as well as they lack of safety, proximity maintenance, convergence guarantees, and robustness to tracking errors and uncertainties. In [10], the authors present a distributed solution, which in theory does not rely on communication and accounts for uncertainties. However, in the experiments they rely on motion capture measurements adding limited noise errors, which is reasonable only in structured environment (i.e., laboratory conditions). In Table  I, we report other state-of-the-art algorithms. Notably, all of them share limitations related to the use of a communication network and/or to not rigorously account for safety, convergence, proximity maintenance and the uncertainties. Conversely in our work, we provide a method which implements all the fore-mentioned features as well as a clear strategy for the selection of the relevant parameters, which in the state-of-the-art is often an unclear or empirical process.
@@ I-B Contribution and paper organization
Our contribution is fivefold:
  • We propose a novel algorithm for flocking control in dense cluttered environments exploiting some of the theoretical findings in [2]. With respect to [3], where the authors introduced the flocking behavior in the Lloyd-based framework, we do not impose the line-of-sight constraint, i.e., two robots are considered in the proximity if they are close enough. This allows to increase the robots mobility and also to not rely on communication. At the same time, it introduces high uncertainties in the position estimation of the neighbors when robots are not in the line of sight.
  • We provide sufficient conditions for safety, proximity maintenance, and convergence even in cluttered environments and with flocking constraints.
  • We propose a novel adaptive method to account for tracking errors and measurement uncertainties.
  • We tailor the use of Convex Weighted Voronoi Diagrams (CWVD) [13] for our problem. By doing that, we considerably improve the performance of the algorithm in dense cluttered environments such as a dense forest.
  • We validate the algorithm in the forest, relying only on on-board sensors i.e., IMU, 2-D Lidar and UVDAR system, without relying neither on a communication network nor GNSS signal.
The paper is organized as follows: Sec. II provides the problem description, in Sec. III we present the background material and the proposed approach, in Sec. IV we provide simulation results and a comparison with a state-of-the-art approach. In Sec. V we validate the proposed approach through experimental results in the field, and finally in Sec. VI we draw the conclusions and the future research directions.
@@ II Problem description
Let us denote the positions of the robots with $p=[p_{1},\dots,p_{N}]^{\top}$, where $p_{i}=[x_{i},y_{i}]^{\top}$ and the obstacle positions by $o=[o_{1},\dots,o_{N_{o}}]^{\top}$, $o_{k}=[x_{k},y_{k}]^{\top}$. The mission space is represented by $\mathcal{Q}$, the weighting function that assess the weights of the points $q\in\mathcal{Q}$ is $\varphi_{i}(q):\mathcal{Q}\rightarrow\mathbb{R}_{+}$, while $\delta_{i}$ and $\delta_{k}^{o}$ are the radius of encumbrance of the $i$–th robot and the $k$–th obstacle respectively.
The problem that we aim to solve reads as follow:
@@
Let $N$ be the number of robots in the swarm, let $\mathcal{N}_{i}$ be the set of neighbors of the robot $i$, and $\bar{\mathcal{N}}_{i}\subseteq\mathcal{N}_{i}$ the set of neighbors that has to remain connected the entire mission. We want to synthesize a distributed control law, such that the swarm maintains the desired proximity constraints, which is characterized by a set of maximum distances $\Gamma_{ij},\,\,\,\forall i,\forall j\in\bar{\mathcal{N}}_{i}$. Each robot in the swarm has to reach its goal region avoiding collisions with other robots and obstacles.
Assumption 1.
Each robot $i$ can perceive the relative position of robot (obstacle) $j$ if $j\in\mathcal{N}_{i}$ ($j\in\mathcal{O}_{i}$). Where $j\in\mathcal{N}_{i}$ if $\|p_{i}-p_{j}\|\leq R_{s}$ ($j\in\mathcal{O}_{i}$ if $\|p_{i}-o_{j}\|\leq R_{s}$), with $R_{s}$ indicating the sensing range, equal for all the robots.
Assumption 2.
The flocking constraints are symmetric, i.e., $\|p_{i}-p_{j}\|\leq\Gamma_{ij},\forall j\in\bar{\mathcal{N}}_{i},\forall i=1,\dots,N$, where $\Gamma_{ij}=\Gamma_{ji}<R_{s}$.
Assumption 3.
The size of the neighboring robots and obstacles is known and it can be approximated with circles/cylinders.
Assumption 4.
In our theoretical analysis, we assume that the robots are holonomic agents.
Assumption 5.
From Section III-C, we assume more complex dynamical systems and a more realistic perception system, where the tracking error is bounded as well as the error on the measured distances from neighbors and obstacles.
@@ III Proposed Approach
The presented approach is grounded in the Lloyd algorithm [27]. In the literature, its distributed nature was exploited to develop novel algorithms for several collective behaviors [3, 2, 13, 28, 29].
Let us consider the following cost function
[…]
where $\mathcal{V}_{i}=\{\mathcal{V}_{i}^{p}\cap\mathcal{V}_{i}^{o}\}$ with $\mathcal{V}_{i}^{p}$ being the $i$–th Convex Weighted Voronoi Diagram (CWVD) introduced in [13], which accounts only for neighboring robots positions
@@
While $\mathcal{V}_{i}^{o}$ is the $i$–th CWVD accounting only for neighboring obstacles positions
@@
where $w=[w^{x},w^{y}]^{\top}=q-p_{i}$ and $\alpha_{ij}=\arctan\left(\frac{y_{j}-y_{i}}{x_{j}-x_{i}}\right)$, while $\epsilon_{p(o)}\in[1,2]$, is an arbitrary real number, which regulates the cautiousness in the robot motion (i.e., $\epsilon_{p(o)}=1$ more aggressive behavior, $\epsilon_{p(o)}=2$ more cautelative behavior, notice that the latter coincides with the classic Voronoi diagram).
@@
By exploiting Assumption 4, we can impose $\dot{p}_{i}=u_{i}$. By following the gradient descent of the cost function (1), we obtain the following proportional control law
@@
is the centroid position evaluated over the $\mathcal{V}_{i}$ cell. By imposing (4), assuming a time-invariant $\varphi_{i}(q)$, each robot $i$, $\forall i,1\dots,N$, converges asymptotically to the centroid position $c_{\mathcal{V}_{i}}$ [30]. In the following, we show how, by introducing suitable modifications to the geometry of cell $\mathcal{V}_{i}$ and crafting an appropriate weighting function $\varphi_{i}(q)$, we can attain a safe and effective algorithm for multi-robot swarming.
III-A Cell geometry $\mathcal{F}_{i}$
The reshape of the cell geometry is both necessary and sufficient to ensure safety guarantees. Hence, we define the cell $\mathcal{A}_{i}=\{\tilde{\mathcal{V}}_{i}^{p}\cap\tilde{\mathcal{V}}_{i}^{o}\cap\mathcal{S}_{i}\}$, where
$\tilde{\mathcal{V}}_{i}^{p}=\left\{\begin{array}[]{l}\left\{q\in\mathcal{V}_{i}^{p}\right\}\text{ if }\frac{1}{2}\left\|p_{i}-p_{j}\right\|\geq\Delta_{ij}\\ \left\{q\in\mathcal{Q}\mid\left\|q-p_{i}\right\|\leq\left\|q-\tilde{p}_{j}\right\|\right\}\text{ otherwise, }\end{array}\right.$ (6)
with $j\in\mathcal{N}_{i}$, $\Delta_{ij}=\delta_{j}+\delta_{i}$, and $\tilde{p}_{j}=p_{j}+2\left(\Delta_{ij}-\frac{\left\|p_{i}-p_{j}\right\|}{2}\right)\frac{p_{i}-p_{j}}{\left\|p_{i}-p_{j}\right\|}$, takes into consideration the physical dimension of the robots. On the other hand,
$\tilde{\mathcal{V}}_{i}^{o}=\left\{\begin{array}[]{l}\left\{q\in\mathcal{V}_{i}^{o}\right\}\text{ if }\frac{1}{2}\left\|p_{i}-p_{j}\right\|\geq\Delta^{o}_{ij}\\ \left\{q\in\mathcal{Q}\mid\left\|q-p_{i}\right\|\leq\left\|q-\tilde{o}_{j}\right\|\right\}\text{ otherwise, }\end{array}\right.$ (7)
with $j\in\mathcal{O}_{i}$, $\Delta^{o}_{ij}=\delta_{i}+\delta^{o}_{j}$ and $\tilde{o}_{j}=o_{j}+2\left(\Delta^{o}_{ij}-\frac{\left\|p_{i}-o_{j}\right\|}{2}\right)\frac{p_{i}-o_{j}}{\left\|p_{i}-o_{j}\right\|}$, takes into consideration the size of the robot and the obstacles, while
@@ III-A Cell geometry $\mathcal{F}_{i}$
with $r_{s,i}$ is half of the sensing radius $R_{s}$.
Theorem 1 (Safety [2]).
By imposing the control $\dot{p}_{i}(\mathcal{A}_{i})$ in (4), collision avoidance is guaranteed at every instant of time.
@@
The proof can be found in [2], the addition of static obstacles and CVWD does not change the result. In particular, by adding static obstacles and/or introducing CVWD does not alter the convexity of the set. ∎
We further apply an additional reshaping of the cell geometry in order to maintain the proximity, i.e., to obtain swarming. By imposing our final cell geometry $\mathcal{F}_{i}=\{\mathcal{A}_{i}\cap\mathcal{M}_{i}\}$, where
$\mathcal{M}_{i}=\{q\in\mathcal{Q}\mid\|q-p_{j}\|<\Gamma_{ij},\forall j\in\bar{\mathcal{N}}_{i}\},$ (9)
we guarantee proximity maintenance while preserving safety. In Figure 3, we depict an example of cell geometry $\mathcal{F}_{i}$.
@@
Given that $\mathcal{F}_{i}\subseteq\mathcal{A}_{i}$, safety comes from Theorem 1. Since $\mathcal{F}_{i}$ is convex by definition, then the centroid $c_{\mathcal{F}_{i}}\in\mathcal{F}_{i}$. It implies that by imposing $\dot{p}_{i}(\mathcal{F}_{i})$ in (4), $p_{i}\in\mathcal{F}_{i}$ at any time. Hence, by the definition of the $\mathcal{F}_{i}$ set, proximity maintenance between $i$ and $j$ is always preserved. ∎

### v2

# Distributed Lloyd-Based Algorithm for Uncertainty-Aware Multi-Robot Under-Canopy Flocking

@@ Abstract
In this letter, we present a distributed algorithm for flocking in complex environments that operates at constant altitude, without explicit communication, no a priori information about the environment, and by using only on-board sensing and computation capabilities. We provide sufficient conditions to guarantee that each robot reaches its goal region in a finite time, avoiding collisions with obstacles and other robots without exceeding a desired maximum distance from a predefined set of neighbors (flocking or proximity constraint). The proposed approach allows to operate in crowded scenarios and to deal with tracking errors and on-board sensing errors, without violating safety and proximity constraints. The algorithm was verified through simulations with varying number of UAVs and also through numerous real-world experiments in a dense forest involving up to four UAVs.
@@ I Introduction
Over the past few decades, numerous applications have emerged for multi-robot systems, exhibiting their undisputed benefits in diverse fields. These applications span domains such as logistics, surveillance, environment monitoring, disaster response, search and rescue, and agriculture among others, demonstrating their versatility. For many of these applications, having a swarm of robots able to flock together towards a desired location can be highly beneficial, due to redundancies and parallelization. The definition of flocking is provided by the seminal work of Reynolds [1], which defines it as a non-colliding aggregate motion. Depending on the environments where robots operate, the reliability or availability of explicit communication between robots may be compromised.
In this work, we propose a Lloyd-based algorithm that relies only on on-board sensors. We aim to synthesize a distributed algorithm, where each robot safely reaches the desired position, while maintaining proximity between robots in the team. We provide sufficient conditions to guarantee safety, proximity maintenance and convergence towards the goal in cluttered environments by accounting for bounded tracking errors and uncertainties in the estimation of neighbors and obstacles’ positions. Our algorithm takes inspiration from the nature; just like flock of birds or a school of fish, the robots do not need a leader and each robot makes its own decisions based on relative information, without relying on explicit communication. We validate our approach through simulations and experiments in the forest (see Figure 1).
Figure 1: Deployment of the proposed algorithm in a real-world scenario with 33 UAVs navigating in a forest.
@@ I-A Related works
Prior literature provides a considerable amount of algorithms for distributed motion coordination of multiple robots, but only a few of them can manage to work in a cluttered environment by also imposing the flocking constraint. Among them, a smaller subset is devoted to applications in the field without relying on explicit communication between the robots.
To position our algorithm within the state of the art, we consider the following features. Distributed control: each robot makes decisions on the basis of local information, which is obtained by relying only on on-board sensors. Asynchronous control: each robot makes decisions independently, in asynchronous fashion. Without explicit communication: the robots do not exchange time-dependent information such as future trajectories, current positions or velocities. Moreover, they do not rely on a common reference frame. Safety: collision and obstacle avoidance is guaranteed at any time. Convergence: the sufficient conditions for the robots to converge to their goal locations are provided. Proximity maintenance: the distance between robots is guaranteed to be below the set threshold on maximum distance. Uncertainties: the system is provably safe and respects proximity constraints even in the presence of bounded tracking errors and/or bounded measurement uncertainties. Cluttered and highly dense environments: the algorithm can operate in cluttered environments such as a forest. In particular, in absence of uncertainties and convergence guarantees, the robots can operate in environments where the minimum distance between two obstacles is large enough to allow the single robot to pass. To ensure convergence and account for uncertainties, a sufficient condition is that the obstacles are spaced far enough to allow passage of twice the robot’s size. On-board localization: the algorithm is verified in unknown environments relying only on on-board localization. Experiments: the algorithm is tested on real robotic platforms in unstructured environments. Open source code: the algorithm is open source.
@@
Table I: List of state-of-the-art algorithms. We indicate with $\times$” if the algorithm does not implement the feature as described in Section I-A, “✓” if it does, and “o” if it does only partially.
In Table I, we report an overview of the most closely related algorithms in the literature. We mark with $\times$” if the algorithm does not implement the feature, “✓” if it does, and finally, “o” if it does but only partially. As shown in Table I, our algorithm is unmatched in fulfilling all the specified features of interest. In fact, despite [5, 4] proposing valid solutions to the problem, their main limitations are the requirement of a trajectory broadcasting network, the absence of both convergence analysis and a rigorous account of tracking errors and uncertainties. Similarly, in [6] the robots need to broadcast information such as poses, velocities, poses covariance, and global extrinsic transformations. On the other hand, [7, 8], do not rely on an explicit communication network, but they exhibit degraded performance in crowded scenarios, lack of safety, proximity maintenance, convergence guarantees analysis, and robustness to tracking errors and uncertainties. In [10], the authors present a distributed solution, which in theory does not rely on explicit communication and accounts for uncertainties. However, in the experiments they rely on motion capture measurements adding limited noise errors, which is reasonable only in structured environment (i.e., laboratory conditions). With respect to [3, 27], where the authors introduced the flocking behavior in the Lloyd-based framework, we do not impose the line-of-sight constraints, i.e., two robots are considered in the proximity if they are close enough. This allows to increase the robots mobility and also to not rely on explicit communication. In Table  I, we report additional relevant state-of-the-art algorithms. Notably, all of them share limitations related to the use of an explicit communication network and/or to not rigorously account for safety, convergence, proximity maintenance and the uncertainties. Conversely, in our work, we provide a method which implements all the fore-mentioned features as well as a comprehensive rule for the selection of the algorithm’s parameters based on measurable quantities (e.g., tracking error, or covariance of state estimates), which is often an unclear or empirical process in the state-of-the-art algorithms.
@@ I-B Contribution and paper organization
Our contribution is threefold. First, we propose a novel algorithm for flocking control in dense cluttered environments. The key novelties lie in our use of an adaptive control strategy to compensate for tracking errors and measurement uncertainties, combined with Convex Weighted Voronoi Diagrams (CWVD) to enhance the performance in a dense forest scenarios. Second, we provide sufficient conditions for safety, proximity maintenance, and convergence in cluttered environments and in the presence of flocking constraints. Finally, we validate the algorithm in the forest, relying only on on-board sensors, i.e., Inertial Measurement Unit (IMU), 2-D Lidar and UVDAR system [28], without relying neither on an explicit communication network nor GNSS signal.
@@ II Problem description
Let us denote the positions of the robots with $p=[p_{1},\dots,p_{N}]^{\top}$, where $p_{i}=[x_{i},y_{i}]^{\top}$ and the obstacle positions by $o=[o_{1},\dots,o_{N_{o}}]^{\top}$, $o_{k}=[x_{k},y_{k}]^{\top}$. The mission space is represented by $\mathcal{Q}\in\mathbb{R}^{2}$, the weighting function that assesses the weights of the points $q\in\mathcal{Q}$ is $\varphi_{i}(q):\mathcal{Q}\rightarrow\mathbb{R}_{+}$, while $\delta_{i}$ and $\delta_{k}^{o}$ are the occupancy radius of the ii–th robot and the kk–th obstacle respectively. The problem that we aim to solve reads as follow:
@@
Let $N$ be the number of robots in the swarm, let $\mathcal{N}_{i}$ be the set of neighbors of the robot $i$, and $\bar{\mathcal{N}}_{i}\subseteq\mathcal{N}_{i}$ the set of neighbors that has to remain connected the entire mission. We want to synthesize a distributed control law, such that the swarm maintains the desired proximity constraints, which is characterized by a set of maximum distances $\Gamma_{ij}=\Gamma_{ji},\,\,\,\forall i\in{1,\dots,N},\forall j\in\bar{\mathcal{N}}_{i}$. Each robot in the swarm has to reach its goal region avoiding collisions with other robots and obstacles.
@@ III Proposed Approach
The presented approach is grounded in the Lloyd algorithm [29]. Let us consider the following cost function
$J_{\operatorname{cov}}({p},\mathcal{V})=\sum_{i\in\mathcal{N}_{i}}\int_{\mathcal{V}_{i}}\left\|q-p_{i}\right\|^{2}\varphi_{i}(q)dq,$ (1)
where $\mathcal{V}=\{\mathcal{V}_{1},\dots,\mathcal{V}_{N}\}$, $\mathcal{V}_{i}=\mathcal{V}_{i}^{p}\cap\mathcal{V}_{i}^{o}$ with $\mathcal{V}_{i}^{p}$ being the ii–th Convex Weighted Voronoi Diagram (CWVD) introduced in [13], which accounts only for neighboring robots’ positions
@@
While $\mathcal{V}_{i}^{o}$ is the ii–th CWVD accounting only for neighboring obstacles positions
@@
where $w_{i}=[w_{i}^{x},w_{i}^{y}]^{\top}=q-p_{i}$ and $\alpha_{ij}=\arctan\left(\frac{y_{j}-y_{i}}{x_{j}-x_{i}}\right)$, while $\epsilon_{p(o)}\in[1,2]$, is a parameter, which regulates the cautiousness in the robot motion (i.e., $\epsilon_{p(o)}=1$ more aggressive behavior, $\epsilon_{p(o)}=2$ more conservative behavior, notice that the latter coincides with the classic Voronoi diagram).
@@
By following the gradient descent of the cost function (1), we obtain the following proportional control law
@@
$c_{\mathcal{V}_{i}}=\frac{\int_{\mathcal{V}_{i}}q\varphi_{i}(q)dq}{\int_{\mathcal{V}_{i}}\varphi_{i}(q)dq}$ (5)
is the centroid position evaluated over the $\mathcal{V}_{i}$ cell. By imposing (4), assuming a time-invariant $\varphi_{i}(q)$, each robot $i\in 1,\dots,N$, converges asymptotically to the centroid position $c_{\mathcal{V}_{i}}$ [30]. In the following, we show how we can attain a safe and effective algorithm for multi-robot swarming by introducing suitable modifications to the geometry of cell $\mathcal{V}_{i}$ and crafting an appropriate weighting function $\varphi_{i}(q)$
III-A Reshaping the cell geometry $\mathcal{V}_{i}$
The reshape of the cell geometry is both necessary and sufficient to ensure safety guarantees. Hence, we define the cell $\mathcal{A}_{i}=\tilde{\mathcal{V}}_{i}^{p}\cap\tilde{\mathcal{V}}_{i}^{o}\cap\mathcal{S}_{i}$, where
@@ III-A Reshaping the cell geometry $\mathcal{V}_{i}$
where $\Delta_{ij}=\delta_{j}+\delta_{i}$, and $\tilde{p}_{j}=p_{j}+2\left(\Delta_{ij}-\frac{\left\|p_{i}-p_{j}\right\|}{2}\right)\frac{p_{i}-p_{j}}{\left\|p_{i}-p_{j}\right\|}$, takes into consideration the physical dimension of the robots. To account for the obstacles,
[…]
where $\Delta^{o}_{ij}=\delta_{i}+\delta^{o}_{j}$ and $\tilde{o}_{j}=o_{j}+2\left(\Delta^{o}_{ij}-\frac{\left\|p_{i}-o_{j}\right\|}{2}\right)\frac{p_{i}-o_{j}}{\left\|p_{i}-o_{j}\right\|}$, takes into consideration the size of the robot and the obstacles, while
@@
where $r_{s,i}$ is half of the sensing radius $R_{s}$ of the robot $i$.
Theorem 1 (Safety).
By imposing the control $\dot{p}_{i}(\mathcal{A}_{i})$ in (4), with $\tilde{\mathcal{V}}_{i}^{p}\cap\tilde{\mathcal{V}}_{i}^{o}\cap\mathcal{S}_{i}$, collision avoidance is guaranteed at every instant of time.
@@
The proof comes from the definition of convex set. Due to the fact that the set $\mathcal{A}_{i}$ is always convex, and by definition $p_{i},c_{\mathcal{A}_{i}}\in\mathcal{A}_{i}$, it follows that since the straight path from the robot position to the centroid position is a safe path, then applying $\dot{p}_{i}(\mathcal{A}_{i})$ in (4) preserves safety. ∎
We further apply an additional reshaping of the cell geometry in order to maintain the desired proximity of the agents, i.e., to achieve swarming behavior. By imposing our final cell geometry $\mathcal{F}_{i}=\mathcal{A}_{i}\cap\mathcal{M}_{i}$, where
$\mathcal{M}_{i}=\{q\in\mathcal{Q}\mid\|q-p_{j}\|\leq\Gamma_{ij},\forall j\in\bar{\mathcal{N}}_{i}\},$ (9)
we obtain proximity maintenance while preserving safety. In Figure 3, we depict an example of cell geometry $\mathcal{F}_{i}$.
@@
Given that $\mathcal{F}_{i}\subseteq\mathcal{A}_{i}$, safety comes from Theorem 1. Since $\mathcal{M}_{i}$ is convex by definition, $\mathcal{F}_{i}$ is also convex, and thus the centroid $c_{\mathcal{F}_{i}}\in\mathcal{F}_{i}$. It implies that by imposing $\dot{p}_{i}(\mathcal{F}_{i})$ in (4), $p_{i}\in\mathcal{F}_{i}$ at any time. Hence, by the definition of the $\mathcal{F}_{i}$ set, proximity maintenance between $i$ and $j$ is always preserved. ∎